Solve 3 / 4 – 1 / 3 Solve 3 / 4 – 1 / 3 Solve 3 / 4 – 1 / 3

Solve
 3 / 4 – 1 / 3
Solve
 3 / 4 – 1 / 3
Solve
 3 / 4 – 1 / 3

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

3 / 4–1 / 3
Taking LCM 12
= [(3 × 3)–(1 × 4)] / 12
= (9–4) / 12
= 5 / 12

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
3/4 - 1/3
Taking LCM as 12,
(3×3)-(1×4)/12
= (9-4)/12
= 5/12
 
Hope it helps!
Thank you ✌️